### v2

@@ 2 Dataset Description and summary statistics
  • Data on transactions submitted via two private mempools, MEV Blocker and Flashbots Protect, obtained from https://Dune.com. [^8] Private mempools are services that allow users to send transactions directly to one or more builders without broadcasting them to the public mempool. At present, there are five major private mempool operators. Four of them—MEV Blocker, Flashbots Protect, Merkel, and Blink—state in their documentation that they share transactions privately with multiple builders. Among these, MEV Blocker and Flashbots Protect publicly disclose which transactions included in a winning block were routed through their systems. The fifth operator, Metamask Smart Transaction, does not disclose its sharing policy.
@@ 2.2 Transactions
The first observation is that the set of proposed transactions is very different from the subset of these transactions that appear in winning blocks. The vast majority (87.7%) of the transactions in our primary dataset that appear in winning blocks are users’ transactions, and only a few are searchers’ transactions. [^11] We classify an address as a searcher if it trades on decentralized exchanges (DEXes) using a smart contract whose source code is not disclosed on https://Etherscan.com. In other words, part of the searcher’s logic is embedded in a smart contract but hidden by making only its bytecode available (the bytecode is the compiled, machine-readable version of the smart contract). Swapping on a DEX is necessary for all common types of “searching,” such as sandwich attacks, arbitrage, backrunning, and triggering liquidations on lending protocols. Also, only few (14.4%) of transactions included in winning blocks are swaps. Instead, among the set of proposed transactions, about half are swaps, which are highly concentrated among a few searchers: the top four addresses alone account for 3,653 swap transactions, or approximately 70% of the total swaps.
Users and searchers submit their transactions to builders in different ways. The default option is to send transactions to the Ethereum public mempool, where builders observe them. Doing so, however, exposes users to attacks. [^12] For example, if a user swaps assets on a blockchain-based financial market (an Automated Market Maker, or AMM), a malicious searcher can execute a “sandwich” attack: it front-runs the user’s swap by making the same trade first, then back-runs it with the opposite swap. This allows the attacker to buy low and sell high, forcing the victim to trade at worse prices. To mitigate this risk, users and searchers may instead use private mempools, services that share transactions privately with multiple builders. Furthermore, some searchers may submit their transactions via direct communication with one or more builders. To be included in a block, a transaction should include either a fee or a payment to the “Fee Recipient” address, a block-level variable set by the builder (typically an Ethereum address they control). The sum of the transfers to the “Fee Recipient” address the fees attached to the transactions in a block is the block’s potential revenues, that is, the paiment that a builder receives if the block is successfully added to the blockchain.
@@
A notable feature of our primary dataset is that 2,981 transactions — 35% of all unique transactions — originate from just two addresses. The first address,1313130x68d3A973E7272EB388022a5C6518d9b2a2e66fBf. is responsible for 1,873 transactions, of which only 31 were included in a winning block. Notably, only 6 (out of 1,873) transactions from this address appear in blocks not built by Titan, and all 6 are simple token transfers to Binance. The second address,1414140x51C72848c68a965f66FA7a88855F9f7784502a7F submitted 1,108 transactions, with just 97 included in a winning block. Only 249 (out of 1,108) transactions appear in blocks not built by Rsync, and only 4 appear in blocks not built by either Rsync or Titan. These 4 transactions are also simple token transfers, including two to addresses associated with Wintermute, an algorithmic trading fund that operates Rsync Builder. Based on these patterns, we refer to them as “Titan-bot” and “Rsync-bot”. [^15] Using a different methodology, Heimbach et al. (2024) also identified Rsync-bot as a non-atomic arbitrage bot, while Titan-bot did not yet exist during the period covered by their study. Wu et al. (2025) also identifies Rsync-bot and Titan-bot as non-atomic arbitrage bots “owned by” or “exclusive to” their respective builders. Almost all transactions from these bots involve token swaps on decentralized exchanges (DEXs): 1,864 for Titan-bot and 1,095 for Rsync-bot. We later discuss how these bots often trade on the same DEX pools during the same auction cycle. Based on this activity, we infer that both bots likely engage in non-atomic arbitrage — that is, arbitrage between on-chain and off-chain financial markets — and are often in competition.
@@
Swaps on decentralized exchanges (DEXes) are also the most common type of transaction across all unique transactions in our primary dataset. We identify 5,291 swap transactions, of which 3,919 occur on Uniswap V3, 1,228 on Uniswap V2, 142 on Sushiswap, and 101 on Pancakeswap. [^16] We observe 99 transactions that interact with multiple DEXes, which is why the total number of swaps on individual DEXes exceeds the number of unique swap transactions. Despite their prevalence, only 814 swap transactions — approximately 15% of all swap transactions — were ultimately included on-chain. A small number of addresses, all belonging to searchers, account for the majority of swaps: the top four addresses are responsible for 3,653 swap transactions (about 70%), with Rsync-bot and Titan-bot alone responsible for 2,959 (56%). Swaps are also concentrated among a few pools: the top 11 DEX pools account for 2,872 transactions (54.4%).
We also identify 807 swap transactions that interact with 42 distinct DEX routers, services that optimize swap execution by splitting trades across multiple liquidity pools and/or using intermediate assets. The most frequently used routers are the Uniswap Universal Router (190 transactions), Uniswap V3 Router (172), 1inch V5 Router (52), and Uniswap V2 Router (42). Of these, 594 transactions were included in a winning block, implying that the majority of swaps included on-chain were routed through a DEX router. The remaining 213 router-based transactions were not included on-chain, and the vast majority of these (153) originate from a single user address, 0xf5213a6a2f0890321712520b8048d9886c1a9900. This user submitted 153 transactions calling the Uniswap V3 Router, of which only two were ultimately included in a winning block.
@@ 3.1 Description of the data for the auction cycle of block 21,322,649
For this auction cycle, our primary dataset contains 669 blocks and 504 unique transactions, of which 300 (59%) are swaps. Only 22 swaps (8 from users and 14 from searchers) were included in winning block 21,322,649, which was selected approximately 13 seconds after the previous winning block. An additional 8 swaps were included in winning block 21,322,650, 3 more in winning block 21,322,651 and 1 more in winning block 21,322,652. Notably, all the swaps included in these later blocks are from users. A total of 181 transactions are attributed to Titan-bot and 42 to Rsync-bot, all of which are DEX swaps. None of the Titan-bot transactions were included in a winning block. In contrast, 5 transactions from Rsync-bot were included, all in winning block 21,322,649.
@@ 3.2 Disparity in Transaction Execution
The last transaction is a user’s swap that remains exclusive to Titan and included in winning block 21,322,652 (also built by Titan). [^20] Its hash is 0x74b6b90aa9ec32adc9e35c15fbd1cf5af37a8ee4d43141785ed21f3bef3f6d8b Hence which builder has access to which transactions affects not only whether transactions are included on-chain, but also when they are included.
@@
Note that all prices here are quoted in ETH. In some cases, such as trades involving stablecoins, we may prefer to express prices in terms of the other token. This simply requires inverting the expressions above.
@@
Finally, we adjust the base currency in the graphs for the implied centralized exchange (CEX) price so that higher prices are always better for the bots. The main result here is that, for the MOG/WETH arbitrage, Titan-bot’s implied CEX price is 1.9% better than that of Rsync-bot (Note that MOG Coin was not listed on Binance nor Coibase during our study period). For USDT/WETH and USDC/WETH, instead, Rsync-bot seems to have an advantage of approximately 0.17% and 0.02%, respectively. Note also that the implied CEX prices are sometimes above or below Binance’s. Remember that the implied CEX prices are risk-adjusted, that is, the arbitrage risk is reflected in a less favorable implied CEX price. Also, the reported Binance price does not account for trading fees or price impact. These reasons may explain why the bots’ implied CEX prices may be below those on Binance. At the same time, the bots may trade on multiple centralized markets or have their own inventory, which they may use to achieve an even better execution price than Binance’s.
@@ 4.1 Time of inclusion in a winning block
We identify 1,288 transactions that were present in multiple bidding cycles. These transactions appeared in a bidding cycle before the winning block was selected, were not included in the winning block, and appeared again in the subsequent bidding cycle. Among these, 871 were present for two cycles, 198 for three cycles, and 108 for four cycles before being included in a winning block. Remarkably, one transaction remained present for 2,437 bidding cycles before inclusion; it was exclusive to the builder payload.de and was ultimately included only when payload.de won. In addition, 28 transactions appeared in multiple cycles but were never included in a winning block. Among the transactions present in multiple bidding cycles, only 17 are searchers’ transactions. Of the remaining transactions, 701 first appeared in our dataset as public, 189 as private, and 381 as exclusive (with 374 exclusive to Titan Builder).
Hence, the vast majority of searchers’ transactions appear in only a single bidding cycle and are either included in the winning block or discarded. Transactions present across multiple bidding cycles are predominantly users’ transactions. Given that our dataset includes 5,947 user transactions, we estimate that approximately 21% users’ transactions are delayed. This delay is partly explained by the fact that 30% of these transactions are, at least initially, exclusive to a single builder who does not win. [^21] For a more in depth analysis, see the companion paper Pahari and Canidio (2025).
@@ 4.2 Successful execution
A transaction may succeed or fail depending on the block in which it is included. To explore this possibility, we consider each transaction hash/block combination and classify it as a fail if its simulation outcome is status = 0, or if it produces no transaction log despite the same transaction generating one in another block. A success is defined as any transaction/block pair not classified as a failure. Under this definition, “failures” include both outright failures and reverts, such as a swap reverting because its execution price would be below the user’s specified worst acceptable price. In contrast, a “success” indicates that the transaction would have executed as intended, had the corresponding block been added to the chain. Among the 49 transactions that sometimes fail and sometimes succeed, 44 are swaps. We therefore focus our analysis exclusively on swaps, which account for 318,742 transaction-block observations.
We create a binary (dummy) variable equal to 1 if a transaction executes successfully, and 0 otherwise. We then run a series of regressions using the following explanatory variables:
@@
We also include transaction fixed effects: that is, a dummy variable for each unique transaction hash. These control for transaction-specific characteristics and capture the average likelihood that a transaction executes successfully across all blocks in which it appears. The coefficients on the other variables then measure how the probability of successful execution varies depending on block timing, builder identity, transaction ordering, and the presence of the two bots.
Finally, we construct two subsamples: swaps in the same direction as those by Titan-bot and Rsync-bot, and swaps in the opposite direction. To do so, for each swap by either Titan-bot or Rsync-bot, we record the target pool, its direction, and the auction cycle (noting that, in our data, all bots’ swaps on a given pool and during the same cycle are in the same direction). We then identify all other swaps occurring in those same pools and cycles, and divide them between those “in the same direction” as the bots and those “in the opposite direction” as the bots.
@@ 4.3 Execution price
For each swap, we compute its average execution price across all blocks, denoted by $p_{\text{avg}}$. We then calculate, for each transaction-block observation, the percentage deviation from this average:
$p_{\text{norm}}=\frac{p-p_{\text{avg}}}{p_{\text{avg}}}\times 100$
where $p$ is the execution price of the swap in a given block. We adjust the base currency so that higher values of $p_{\text{norm}}$ always reflect better execution from the user’s perspective. The distribution of $p_{\text{norm}}$ has a median of $1.41\times 10^{-14}$, a minimum of 10.53-10.53 and a maximum of $13.47$ basis points.
Table 2 reports the coefficients from three regressions of $p_{\text{norm}}$ on the same set of explanatory variables and subsamples used in the success/failure analysis. As with the previous table, the regression on the full sample yields ambiguous results: being included in a block built by Titan is associated with slightly worse execution prices, while the presence of a transaction from Titan-bot in the same block improves execution. For Rsync, the pattern is reversed. These opposing effects largely offset each other. In contrast, the regressions restricted to swaps in the same direction as either bot show a clearer trend: execution prices are between 9 and 12 basis points worse when such swaps are included in blocks built by Rsync or Titan (the presence of the two bots has a negligible effect). For swaps in the opposite direction from the bots, results align with those from the success/failure regressions: being included in a block built by Rsync or Titan improves the execution price by between 9 and 27 basis points. The effect remains positive (but smaller) if the block also contains a Rsync-bot or Titan-bot transaction. Additionally, being placed later in the block (i.e., with a higher transaction index) has a small but consistently positive effect on price. Finally, while many coefficients are statistically significant, the regressions explain only a modest share of the variation in $p_{\text{norm}}$, as reflected in the low $R^{2}$ values.
Table 2: Normalized price $p_{norm}$

### v3

@@ 2 Dataset Description and summary statistics
  • Data on transactions submitted via two private mempools, MEV Blocker and Flashbots Protect, obtained from https://Dune.com.
@@ 2.2 Transactions
The first observation is that the set of proposed transactions is very different from the subset of these transactions that appear in winning blocks. The vast majority (87.7%) of the transactions in our primary dataset that appear in winning blocks are users’ transactions, and only a few are searchers’ transactions. [^11] We classify an address as a searcher if it trades on decentralized exchanges (DEXes) using a smart contract whose source code is not disclosed on https://Etherscan.com. In other words, part of the searcher’s logic is embedded in a smart contract but hidden by making only its bytecode available (the bytecode is the compiled, machine-readable version of the smart contract). Swapping on a DEX is necessary for all common types of “searching,” such as sandwich attacks, arbitrage, backrunning, and triggering liquidations on lending protocols. Also, only few (14.4%) of transactions included in winning blocks are swaps. Instead, among the set of proposed transactions, about half are swaps. Also, about half of proposed transactions are from searchers.
To be included in a block, a transaction should include either a fee or a payment to the “Fee Recipient” address, a block-level variable set by the builder (typically an Ethereum address they control). The sum of the transfers to the “Fee Recipient” address the fees attached to the transactions in a block is the block’s potential revenues, that is, the paiment that a builder receives if the block is successfully added to the blockchain.
@@
A notable feature of our primary dataset is that 2,981 transactions — 35% of all unique transactions — originate from just two addresses. The first address1212120x68d3A973E7272EB388022a5C6518d9b2a2e66fBf. is responsible for 1,873 transactions, of which only 31 were included in a winning block. Notably, only 6 (out of 1,873) transactions from this address appear in blocks not built by Titan, and all 6 are simple token transfers to Binance. The second address1313130x51C72848c68a965f66FA7a88855F9f7784502a7F submitted 1,108 transactions, with just 97 included in a winning block. Only 249 (out of 1,108) transactions appear in blocks not built by Rsync, and only 4 appear in blocks not built by either Rsync or Titan. These 4 transactions are also simple token transfers, including two to addresses associated with Wintermute, an algorithmic trading fund that operates Rsync Builder. Based on these patterns, we refer to them as “Titan-bot” and “Rsync-bot”. [^14] Using a different methodology, heimbach2024non also identified Rsync-bot as a non-atomic arbitrage bot, while Titan-bot did not yet exist during the period covered by their study. wu2025measuring also identifies Rsync-bot and Titan-bot as non-atomic arbitrage bots “owned by” or “exclusive to” their respective builders. Almost all transactions from these bots involve token swaps on decentralized exchanges (DEXs): 1,864 for Titan-bot and 1,095 for Rsync-bot. We later show that these bots frequently trade on the same DEX pools within the same auction cycle. Moreover, when they trade the same token pairs in the same direction, their swaps are almost always mutually exclusive, in the sense that no proposed block contains transactions from both bots. [^15] There is a single exception to this pattern: a Titan-bot transaction selling approximately USD 700 of MOG Coin for ETH appears in some blocks alongside an Rsync-bot transaction selling approximately USD 7,000 of MOG Coin for ETH. Taken together, this trading behavior indicates that both bots primarily engage in non-atomic arbitrage—that is, arbitrage between on-chain and off-chain markets—and frequently compete for the same opportunities.
@@
Swaps on decentralized exchanges (DEXes) are also the most common type of transaction across all unique transactions in our primary dataset. We identify 5,282 swap transactions, of which 3,915 occur on Uniswap V3, 1,222 on Uniswap V2, 142 on Sushiswap, and 102 on Pancakeswap. [^16] We observe 99 transactions that interact with multiple DEXes, which is why the total number of swaps on individual DEXes exceeds the number of unique swap transactions. Despite their prevalence, only 808 swap transactions — approximately 15% of all swap transactions — were ultimately included on-chain. A small number of addresses, all belonging to searchers, account for the majority of swaps: the top four addresses are responsible for 3,649 swap transactions (about 70%), with Rsync-bot and Titan-bot alone responsible for 2,955 (56%). Swaps are also concentrated among a few pools: the top 11 DEX pools account for 2,876 transactions (54.4%).
We also identify 794 swap transactions that interact with 41 distinct DEX routers, services that optimize swap execution by splitting trades across multiple liquidity pools and/or using intermediate assets. The most frequently used routers are the Uniswap Universal Router (190 transactions), Uniswap V3 Router (172), 1inch V5 Router (52), and Uniswap V2 Router (42). Of these, 581 transactions were included in a winning block, implying that the majority of swaps included on-chain were routed through a DEX router. The remaining 213 router-based transactions were not included on-chain, and the vast majority of these (151) originate from a single user address, 0xf5213a6a2f0890321712520b8048d9886c1a9900. This user submitted 153 transactions calling the Uniswap V3 Router, of which only two were ultimately included in a winning block.
@@ 3.1 Description of the data for the auction cycle of block 21,322,649
For this auction cycle, our primary dataset contains 669 blocks and 504 unique transactions, of which 299 (59%) are swaps. Only 21 swaps (8 from users and 13 from searchers) were included in winning block 21,322,649, which was selected approximately 13 seconds after the previous winning block. An additional 8 swaps were included in winning block 21,322,650, 3 more in winning block 21,322,651 and 1 more in winning block 21,322,652. Notably, all the swaps included in these later blocks are from users. A total of 181 transactions are from Titan-bot and 42 are from Rsync-bot, all of which are DEX swaps. None of the Titan-bot transactions were included in a winning block. In contrast, 5 transactions from Rsync-bot were included, all in the winning block 21,322,649.
@@ 3.2 Disparity in Transaction Execution
The last transaction is a user’s swap that remains exclusive to Titan and is finally included in winning block 21,322,652 (also built by Titan). [^20] Its hash is 0x74b6b90aa9ec32adc9e35c15fbd1cf5af37a8ee4d43141785ed21f3bef3f6d8b Hence which builder has access to which transactions affects not only whether transactions are included on-chain, but also when they are included.
@@
Note that all prices here are quoted in ETH. In some cases, such as trades involving stablecoins, we may prefer to express prices in terms of the other token. This simply requires inverting the expressions above. Finally, each observation of $p_{\text{implied}}$ has a timestamp equal to the timestamp of the block including the swap transaction representing the DEX leg of the trade.
@@
Finally, in both Figure 5 and 6, we adjust the base currency in the graphs for the implied centralized exchange (CEX) price so that higher prices are better for the bots. The main result here is that, for the MOG/WETH arbitrage, Titan-bot’s implied CEX price is 1.9% better than that of Rsync-bot (Note that MOG Coin was not listed on Binance nor Coibase during our study period). For USDT/WETH and USDC/WETH, instead, Rsync-bot seems to have an advantage of approximately 0.17% and 0.02%, respectively. Note also that the implied CEX prices are sometimes above or below Binance’s. Remember that the implied CEX prices are risk-adjusted, that is, the arbitrage risk is reflected in a less favorable implied CEX price. Also, the reported Binance price does not account for trading fees or price impact. These reasons may explain why the bots’ implied CEX prices may be below those on Binance. At the same time, the bots may trade on multiple centralized markets or have their own inventory, which they may use to achieve an even better execution price than Binance’s.
@@ 4.1 Time of inclusion in a winning block
We identify 1,288 transactions that were present in multiple bidding cycles. These transactions appeared in a bidding cycle before the winning block was selected, were not included in the winning block, and appeared again in the subsequent bidding cycle. Among these, 871 were present for two cycles, 198 for three cycles, and 108 for four cycles before being included in a winning block (see Figure 7). Remarkably, one transaction remained present for 2,437 bidding cycles before inclusion; it was exclusive to the builder payload.de and was ultimately included only when payload.de won. In addition, 28 transactions appeared in multiple cycles but were never included in a winning block. Among the transactions present in multiple bidding cycles, only 17 are searchers’ transactions. Of the remaining transactions, 701 first appeared in our dataset as public (i.e., present also in the public mempool), 189 as private (i.e., present in blocks by multiple builders but not in the public mempool), and 381 as exclusive to a single builder (with 374 exclusive to Titan Builder).
Figure 7: Empirical CDF of the waiting time to on-chain inclusion, measured in blocks, for users’ transactions.
Hence, the vast majority of searchers’ transactions appear in only a single bidding cycle and are either included in the winning block or discarded. Transactions present across multiple bidding cycles are predominantly users’ transactions. Given that our dataset includes 5,968 user transactions, we estimate that approximately 21% users’ transactions are delayed. This delay is partly explained by the fact that 30% of these transactions are, at least initially, exclusive to a single builder who does not win.
In addition, 12 transactions that were delayed because exclusive eventually appear in the public mempool. Nine of these were exclusive for two cycles and three for three cycles. In four cases, after appearing in the public mempool, these transactions were added on chain by the same builder that had them as exclusive to start with (Titan). All appear to be from users: three used the 0x router and four the 1inch router. The existence of these transactions is problematic, as they were first delayed and later exposed in the public mempool, leaving users vulnerable to attacks (and one of them appears to have been sandwiched). [^21] These transactions are: 0x1daea1584d684385fb25209bad3a49c54a1e2c500e27543334c7b3e695a12ffb, 0x6605c1c36fcd3d696844519d6ff7199fe4ddd34fb4f3e0ed4da71aedfcf1c633, 0x3dadbab8e85b6e48a2c33f9029ea4ebebdd3def8b9d907faff15ef81527de968, 0x9c7364eddb94fe659bc8104585ff86a8a234480fe00a9f9aa3eb410e572ebdf2, 0x9d476dbc524651cbe26bcafab99f6fe40a823a5588fa5cd5002e5b414a8a2f8c, 0x417440b083bdcb529663c6615c32d89a92beb7863f1a5d6b12bca5b94f4a0f87 (sandwiched according to https://www.zeromev.org/block?num=21322643), 0xe56caee27e79dee8a4c7ed6f945ccbc2cc733d761e19bd68329dc364c61e73a7, 0xae5b88bc0bffa04811eda0dfdcd7224599e2757339e18b2d57f4f40d61902a4c, 0xd41139b67a193045b4d60266c084f6dadd0cb95dccd7b2a06ea68653d6a313ec, 0x597fc0cb0610d1e3ea2baf0f6f90d0d741a6429bfdd9b375bfd2493b70d41dcd, 0x44e8105d5c672195877dc1ed970511e0f0a15fc25cbcafe71d51d428d54c215d, 0xd74718e6d8d58bc33334c0c2f6642d9dec49496329329d8fdcb1e698e81abc94.
Unfortunately, our data do not allow us to determine with certainty which wallet was used to create a given transaction, which private mempool (or other service) was used to relay it to builders, or how the transaction was propagated through the public mempool. This information is simply not recorded in Ethereum blocks. Nonetheless, a plausible explanation is that some private mempool operators share transactions exclusively with a single builder, reverting to the public mempool if that builder fails to win two or three consecutive auctions.
@@ 4.2 Successful execution
A transaction may execute successfully or fail depending on the transactions that precede it. Because a given transaction may be preceded by different sets of transactions across the blocks in which it appears, the same transaction may execute successfully in some proposed blocks but fail in others.
To study this possibility, we consider each transaction–block pair and classify it as a failure if its simulated execution yields status = 0, or if it produces no transaction log despite generating one when included in another block. A success is defined as any transaction–block pair not classified as a failure. Under this definition, failures include both outright execution failures and reverts, such as a swap reverting because its execution price would fall below the user’s specified minimum acceptable price. By contrast, a success indicates that the transaction would have executed as intended had the corresponding block been added to the chain.
We identify 149 transactions that always fail and 49 transactions that succeed in some blocks but fail in others. Of these 49 transactions, 44 are swaps: 10 generated by searchers and 34 by users. Because our dataset contains 10,793 transactions—of which 4,565 are searchers’ swaps and 717 are users’ swaps—this implies that virtually all searcher swaps always execute successfully, while 3.3% of users’ swaps fail in at least one of the blocks in which they appear. Virtually all remaining transactions either always succeed or always fail. [^22] A subtlety arises because transactions that always fail does not generate any log and cannot be classified by type. As a result, identifying a transaction as a swap requires that it execute successfully in at least one block.
We therefore focus on user swaps that sometimes fail and plot the empirical CDF of their fraction of failures (Figure 8). Approximately 50% of these transactions fail in at least half of the candidate blocks in which they appear, and roughly 30% fail in more than 80% of such blocks. The main takeaway is that the probability a user’s swap fails is highly skewed: the vast majority of users’ swaps (96.7%) never fail, while a small fraction fail very frequently—1.7% fail in at least half of the blocks in which they appear, and 1.1% fail in more than 80% of such blocks. This skewness indicates that execution risk is concentrated among a small subset of user swaps, rather than being evenly distributed.
Figure 8: Empirical CDF of the fraction of candidate blocks in which a given user swap fails, conditional on the swap sometimes failing and sometimes succeeding.
We then perform a regression analysis restricted to swaps that both fail and succeed, which account for 318,742 transaction–block observations. For each swap $i$ appearing in block $k$, we construct a binary outcome variable equal to one if transaction $i$ executes successfully in block $k$, and zero otherwise. We then regress this outcome on the following explanatory variables:
@@
We also include transaction fixed effects: that is, a dummy variable for each unique transaction hash. These control for transaction-specific characteristics and capture the average likelihood that a transaction executes successfully across all blocks in which it appears. The coefficients on the other variables then measure how the probability of successful execution varies depending on block timing, builder identity, transaction ordering, and the presence of the two bots. Finally, we construct two subsamples: swaps in the same direction as those by Titan-bot and Rsync-bot, and swaps in the opposite direction. To do so, for each swap by either Titan-bot or Rsync-bot, we record the target pool, its direction, and the auction cycle (noting that, in our data, all bots’ swaps on a given pool and during the same cycle are in the same direction). We then identify all other swaps occurring in those same pools and cycles, and divide them between those “in the same direction” as the bots and those “in the opposite direction” as the bots.
@@ 4.3 Execution price
For each swap $i$, we compute its average execution price across all blocks, denoted by $p_{\text{avg},i}$. We then calculate, for each transaction-block observation, the percentage deviation from this average:
$p_{dev,i,k}\equiv\frac{p_{i,k}-p_{\text{avg},i}}{p_{\text{avg},i}}\times 100$
where $p_{i,k}$ is the execution price of swap $i$ in proposed block $k$. We adjust the base currency so that higher values of $p_{dev,i,k}$ always reflect better execution from the user’s perspective. The distribution of $p_{dev,i,k}$ across all $i$ and $k$ has a median of $1.41\times 10^{-14}$, a minimum of 10.53-10.53 and a maximum of $13.47$ basis points.
Figure 9: The CDF of the standard deviation of the $p_{dev,i,k}$.
For each transaction $i$, we compute the standard deviation of $p_{\text{dev},i,k}$. This yields one observation per swap, capturing the dispersion of its execution price across candidate blocks. Figure 9 plots the distribution of these standard deviations. As before, the distribution is extremely skewed: the median is 0.043, while the mean is 0.27. Moreover, for 15% of swaps the standard deviation of $p_{\text{dev},i,k}$ is below $10^{-9}$, whereas for 7% of swaps it exceeds $1$. For the sake of interpreting these magnitudes, suppose that for each swap $i$ the distribution of $p_{\text{dev},i,k}$ is normal. Under this assumption, for half of the swaps in our sample there is less than a one-third probability that the execution price deviates by more than 4.3 basis points from $p_{\text{avg},i}$, while for 7% of swaps there is more than a one-third probability that the execution price deviates by at least 100 basis points from $p_{\text{avg},i}$.
Also here, we resort to a regression analysis to decompose such variation. Table 2 reports the coefficients from three regressions of $p_{dev,i,k}$ on the same set of explanatory variables and subsamples used in the success/failure analysis. As with the previous table, the regression on the full sample yields ambiguous results: being included in a block built by Titan is associated with slightly worse execution prices, while the presence of a transaction from Titan-bot in the same block improves execution. For Rsync, the pattern is reversed. These opposing effects largely offset each other. In contrast, the regressions restricted to swaps in the same direction as either bot show a clearer trend: execution prices are between 9 and 12 basis points worse when such swaps are included in blocks built by Rsync or Titan (the presence of the two bots has a negligible effect). For swaps in the opposite direction from the bots, results align with those from the success/failure regressions: being included in a block built by Rsync or Titan improves the execution price by between 9 and 27 basis points. The effect remains positive (but smaller) if the block also contains a Rsync-bot or Titan-bot transaction. Additionally, being placed later in the block (i.e., with a higher transaction index) has a small but consistently positive effect on price. Finally, while many coefficients are statistically significant, the regressions explain only a modest share of the variation in $p_{dev,i,k}$, as reflected in the low $R^{2}$ values.
Table 2: Price deviations $p_{dev,i,k}$